undeadknight26 (undeadknight26):

Which of the following tables shows the correct steps to transform x2 + 6x + 8 = 0 into the form (x - p)2 = q?

undeadknight26 (undeadknight26):

I say (x + 3)^2 = -2

OpenStudy (ranga):

x^2 + 6x + 8 = 0 x^2 + 6x = -8 (x+3)^2 = -8 + 9 (x+3)^2 = 1 (x - (-3))^2 = 1

undeadknight26 (undeadknight26):

So it would be: (x + 3)^2 = 1?

OpenStudy (ranga):

Yes. Also they want it in the form (x-p)^2 = q and so (x - (-3))^2 = 1 so that p = -3 and q = 1.

undeadknight26 (undeadknight26):

Thank you good sir!